OpenStudy (anonymous):

Solve for x. |x – 2| + 10 = 12 x = 4 and x = 0 x = -15 and x = 9 x = -4 and x = -2 x = -3 and x = 1 no solutions

OpenStudy (ash2326):

We have \[|x-2|+10=12\] do you know what's |x|?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x-2+10=12>>4 -x+2+10=12>>0

OpenStudy (anonymous):

You have to use X-2+10=12 X=4 And you have to use X+2+10=12 X=0 So it's the first choice

OpenStudy (ash2326):

@moody22 |x| is the absolute value function if x>=0 then |x|=x if x< 0 then |x|=-x

OpenStudy (anonymous):

You can also input the problem into wolframalpha and it gives the same answer I did.

OpenStudy (anonymous):

is my answer is correct @ash2326

OpenStudy (ash2326):

@moha_10 Yeah you are right

OpenStudy (anonymous):

thx